OpenStudy (anonymous):

Solve the system using elimination 3x + 3y = –9 3x – 3y = 21 a. (3, –6) b. (–5, 2) c. (3, 3) d. (2, –5)

OpenStudy (anonymous):

The answer is d. (2,-5)

OpenStudy (anonymous):

The y's cancel out leaving you with \[6x = 12\] Divide by 6 and x=2.. Plug in x =2 into any of the equations \[3(2) + 3y = -9\] Evaluate and solve for y \[6 + 3y = -9\] \[3y = -15\] \[y = -5\] Thus the answer is d.

OpenStudy (anonymous):

3x + 3y = -9 3x - 3y = 21 ______________ Just add them, 6x + 0y = 12 6x = 12 x = 2 Now that you have x, which is 2, plug 2 back in one of the equations. 3(2) + 3y = -9 6 + 3y = -9 -6 -6 Subtract six from both sides. 3y = -15 y = -5 Now you have both x and y, just write them both as coordinates, (2,-5)